# X-ray Variability in the Symbiotic Binary RT Cru: Principal Component Analysis

### v1

@@ 5.2 Flickering Behavior
Binary systems hosting degenerate cores usually manifest flickering, which is associated with accretion physics (see, e.g., Luna et al., 2013; Merc et al., 2024). RT Cru is characterized by flickering, referring to the stochastic variability in the light curve characterized by less than tenths of a magnitude, occurring on scales from seconds to minutes. The flickering behavior in this object has been recorded from the optical band (Cieslinski et al., 1994) to UV (Luna et al., 2018) and X-rays (Ducci et al., 2016; Danehkar et al., 2021). This behavior was further seen in the B, V, and R bands from the ground-based observations and photometry made with the NASA TESS mission (Pujol et al., 2023). Its TESS light curves show accumulation-induced flickering variability on timescales of minutes, with substantial variation in flickering (Merc et al., 2024). Our variability simulations suggest that changes in the column density or/and covering fraction of the absorbing material, along with the X-ray brightness from accretion processes, could potentially lead to long-term variations in the soft excess below 4 keV, which aligns with the spectral fitting results of Luna et al. (2018). Moreover, simulated X-ray spectra indicate that rapid flickering-type variations might be caused by some changes in the thermal plasma emission, which Luna & Sokoloski (2007) found likely originates from a boundary layer of the accretion disk around a massive white dwarf rather than a magnetically channeled flow.
The disappearance of optical Balmer emission lines and decreases in , , and flickering amplitudes were also recorded in RT Cru in 2019, attributed to a decline in the accretion process, followed by its re-appearance in the later years associated with restoring the accretion flow (Pujol et al., 2023). In particular, flickering in the -type symbiotic star T CrB is predominantly detected in hard X-rays during two active phases with periods of d and d, which seems to be produced in the boundary layer due to variable mass transfer (Iłkiewicz et al., 2016). Moreover, flickering in the -class symbiotic star AG Peg, whose soft emission may be a result of CSW, resembles those produced by accretion processes in the X-ray spectra over 2013–2015 (Zhekov & Tomov, 2016). Rapid, low-amplitude UV flickering has been seen in and / sources, which seem to originate from accretion processes rather than quasi-steady thermonuclear burning on the white dwarf surface, or CSW (Luna et al., 2013).
@@ 6 Conclusion
(iv) PCA of the two Suzaku XIS1 observations provides us with the three spectral components: The first PCA spectrum likely corresponds to the line-of-sight absorbing material since the corresponding light curve rises with decreases in the brightness as seen in Fig. 3. The second component is probably associated with an absorbed continuum, consisting of a soft blackbody-like spectrum and a hard spectrum. Based on our simulations of Suzaku spectra, the X-ray variations are mostly caused by changes in the absorbing columns, with some effects also owing to alterations in the continuum. The last PCA spectrum contains some spectral features akin to those seen in of the XMM-Newton data, so it may be related to the heavily obscured, soft plasma emission with keV suggested by Danehkar et al. (2021) using low-count Bayesian statistics. Additionally, a similar collisional plasma temperature is derived from a phenomenological model fitted to the reconstructed spectra of the first and second PCA components of the Suzaku XIS1 data, as well as a simple model matched to the XMM-Newton observation. Such a thermal emission feature can be created with a wind velocity of 1,000 km s-1 similar to jets found in CH Cyg (Karovska et al., 2007, 2010). Our simulations of the XIS1 spectra also illustrate that some changes in the thermal plasma emission could lead to the third PCA component of Suzaku, resulting in flickering-type variations.
@@
In summary, our eigenvector-based multivariate analysis of the Suzaku data of RT Cru suggests that the hardness transition seen over two epochs is primarily related to changes in the absorbing material and partially the X-ray brightness caused by accretion processes. In addition, our analysis of the multi-mission X-ray data taken with different telescopes supports the likelihood of a soft thermal plasma emission component, as previously proposed by Danehkar et al. (2021) with Bayesian analysis, which may be heavily obscured by the line-of-sight absorbing material as well as high levels of background noise. A recent statistical approach by Zhang et al. (2023) could not robustly constrain any soft-band emission lines, but it did provide an upper confidence level of 1 keV for the plasma temperature. Furthermore, changes in thermal plasma emission from an accretion disk boundary layer likely cause the flickering-type variations observed in this object. Some future X-ray telescopes, such as the proposed Arcus (Smith et al., 2016), with a much higher sensitivity below 1 keV, will be able to capture the emission features in the soft band. Moreover, future high-spectral resolution observations with the recently launched telescope XRISM (Tashiro et al., 2020) will certainly help us disclose further details of the X-ray spectral features of RT Cru.

### v2

@@ 5.2 Flickering Behavior
Binary systems hosting degenerate cores usually manifest flickering, which is associated with accretion physics (see, e.g., Luna et al., 2013; Merc et al., 2024). RT Cru is characterized by flickering, referring to the stochastic variability in the light curve characterized by less than tenths of a magnitude, occurring on scales from seconds to minutes. The flickering behavior in this object has been recorded from the optical band (Cieslinski et al., 1994) to UV (Luna et al., 2018) and X-rays (Ducci et al., 2016; Danehkar et al., 2021). This behavior was further seen in the , , and bands from the ground-based observations and photometry made with the NASA TESS mission (Pujol et al., 2023). Its TESS light curves show accumulation-induced flickering variability on timescales of minutes, with substantial variation in flickering (Merc et al., 2024). Our variability simulations suggest that changes in the column density or/and covering fraction of the absorbing material, along with the X-ray brightness from accretion processes, could potentially lead to long-term variations in the soft excess below 4 keV, which aligns with the spectral fitting results of Luna et al. (2018). Moreover, simulated X-ray spectra indicate that rapid flickering-type variations might be caused by some changes in the thermal plasma emission, which Luna & Sokoloski (2007) found likely originates from a boundary layer of the accretion disk around a massive white dwarf rather than a magnetically channeled flow.
The disappearance of optical Balmer emission lines and decreases in , , and flickering amplitudes were also recorded in RT Cru in 2019, attributed to a decline in the accretion process, followed by its reappearance in the later years associated with restoring the accretion flow (Pujol et al., 2023). In particular, flickering in the -type symbiotic star T CrB is predominantly detected in hard X-rays during two active phases with periods of d and d, and seems to be produced in the boundary layer due to variable mass transfer (Iłkiewicz et al., 2016). Moreover, flickering in the -class symbiotic star AG Peg, whose soft emission may be a result of CSWs, resembles those produced by accretion processes in the X-ray spectra over 2013–2015 (Zhekov & Tomov, 2016). Rapid, low-amplitude UV flickering has been seen in and / sources, which seem to originate from accretion processes rather than quasi-steady thermonuclear burning on the white dwarf surface, or CSWs (Luna et al., 2013).
@@ 6 Conclusion
(iv) PCA of the two Suzaku XIS1 observations provides us with the three spectral components. The first PCA spectrum likely corresponds to the line-of-sight absorbing material since the corresponding light curve rises with decreases in the brightness as seen in Fig. 3. The second component is probably associated with an absorbed continuum, consisting of a soft blackbody-like spectrum and a hard spectrum. Based on our simulations of Suzaku spectra, the X-ray variations are mostly caused by changes in the absorbing columns, with some effects also owing to alterations in the continuum. The last PCA spectrum contains some spectral features akin to those seen in of the XMM-Newton data, so it may be related to the heavily obscured, soft plasma emission with keV suggested by Danehkar et al. (2021) using low-count Bayesian statistics. Additionally, a similar collisional plasma temperature is derived from a phenomenological model fitted to the reconstructed spectra of the first and second PCA components of the Suzaku XIS1 data, as well as a simple model matched to the XMM-Newton observation. Such a thermal emission feature can be created with a wind velocity of 1000 km s-1 similar to jets found in CH Cyg (Karovska et al., 2007, 2010). Our simulations of the XIS1 spectra also illustrate that some changes in the thermal plasma emission could lead to the third PCA component of Suzaku, resulting in flickering-type variations.
@@
In summary, our eigenvector-based multivariate analysis of the Suzaku data of RT Cru suggests that the hardness transition seen over two epochs is related primarily to changes in the absorbing material and partially to the X-ray brightness caused by accretion processes. In addition, our analysis of the multi-mission X-ray data taken with different telescopes supports the likelihood of a soft thermal plasma emission component, as previously proposed by Danehkar et al. (2021) with Bayesian analysis, which may be heavily obscured by the line-of-sight absorbing material as well as high levels of background noise. A recent statistical approach by Zhang et al. (2023) could not robustly constrain any soft-band emission lines, but it did provide an upper confidence level of 1 keV for the plasma temperature. Furthermore, changes in thermal plasma emission from an accretion disk boundary layer likely cause the flickering-type variations observed in this object. Some future X-ray telescopes, such as the proposed Arcus (Smith et al., 2016), with a much higher sensitivity below 1 keV, will be able to capture the emission features in the soft band. Moreover, future high-spectral resolution observations with the recently launched telescope XRISM (Tashiro et al., 2020) will certainly help us disclose further details of the X-ray spectral features of RT Cru.